# How Far Can Transformers Reason? The Globality Barrier and Inductive Scratchpad

### v2

@@ Abstract
Can Transformers predict new syllogisms by composing established ones? More generally, what type of targets can be learned by such models from scratch? Recent works show that Transformers can be Turing-complete in terms of expressivity, but this does not address the learnability objective. This paper puts forward the notion of globality degree of a target distribution to capture when weak learning is efficiently achievable by regular Transformers, where the latter measures the least number of tokens required in addition to the tokens histogram to correlate nontrivially with the target. As shown experimentally and theoretically under additional assumptions, distributions with high globality cannot be learned efficiently. In particular, syllogisms cannot be composed on long chains. Furthermore, we show that (i) an agnostic scratchpad cannot help to break the globality barrier, (ii) an educated scratchpad can help if it breaks the globality at each step, however not all such scratchpads can generalize to out-of-distribution (OOD) samples (iii) a notion of ‘inductive scratchpad’, that composes the prior information more efficiently, can both break the globality barrier and improve the OOD generalization. In particular, some inductive scratchpads can achieve length generalizations of up to 6×6$\times$6 × for some arithmetic tasks depending on the input formatting.
@@ 1.1 Syllogisms composition
This type of task is nontrivial for current LLMs, and we refer to Appendix G for some further experiments with GPT models. [^2] At the time of the experiments, ChatGPT was in particular not successful at these two tasks. Note that here we are not interested specifically in solving a graph-based task, but rather in understanding when Transformers can compose and more generally how far they can do so. We would like to identify particular measures on the data distribution (e.g., syllogisms topologies in the above example) that capture when Transformers can efficiently learn.
@@ 1.2 Hardness of long compositions
The input of this task is the graph edge set with the queried vertices. The label is 0 if the two queried vertices are not connected (Class 1) and 1 if they are (Class 2). We refer to this task as the ‘cycle task’. See Figure 1(a) for an illustration.
@@
(b) Nb. of iterations (y-axis) with $512$ fresh samples to learn ($\geq 95\%$ acc.) the cycle task of size $n$ (x-axis) using GPT2-style models with $10M,25M,85M$ parameters.
@@
A challenge for the cycle task is that there is no clear ‘low-complexity pattern’ in the input representation that indicates whether there are 1 or 2 cycles. No simple statistics based on degrees, edge counts, or finite motif counts that can tell if the vertices are connected or not. One has to consider at least $n$ edges in order to get any correlation with the presence of a path. In other words, the task requires a ‘global reasoning’ involving a ‘large’ number of input tokens and this seems to be hard for Transformers.
@@ 1.3 Hardness of global reasoning
As discussed previously, the cycle task appears to be challenging for Transformers as it requires some global reasoning. Other tasks such as subset parities exhibit the same challenge. However the latter can be proved to be not efficiently learnable by various regular neural networks and noisy gradient descent, as one can get explicitly a class of functions (through orbit arguments [12, 13]) that has large statistical dimension [14] or low cross-predictability [12, 15] (see Appendix A.2). For the cycle task, we have a single distribution, and it is unclear how to use the invariances of Transformers to get arguments as in [12, 13], as the input distribution is not invariant under the symmetries of the model. We thus would like to develop a more general complexity measure that unifies why such tasks are hard for Transformer-like models and that formalizes the notion of ‘global reasoning barrier’ when models are trained from scratch. We also would like to understand how the scratchpad methodologies that have proved helpful in various settings (see Section 3) can help here. This raises the questions:
@@ 1.4 Our contributions
    • A theorem (Theorem 1) that proves the negative side of the above conjecture, the globality barrier, in the instance of a variant of the cycle task under certain technical assumptions. (The cycle task is also put forward in the paper as a simple benchmark to test the global reasoning capabilities of models.)
@@
    • Agnostic scratchpad: we extend Theorem 1 to cases where a polynomial-size scratchpad is used by the Transformer, without any supervision of the scratchpad. I.e., the scratchpad gives additional memory space for the Transformer to compute intermediate steps. This shows that efficient weak learning is still not possible with such an agnostic scratchpad if the globality is non-constant. An educated guess about what to learn in the scratchpad based on some target knowledge is thus required.
@@
    • We introduce the notion of inductive scratchpad, a type of educated scratchpad that exploits ‘induction’ compared to a fully educated scratchpad and thus composes more efficiently the prior state information. We show that when the target admits an inductive decomposition, such as for the cycle, arithmetic, or parity tasks, the inductive scratchpad both breaks the globality and improves the OOD generalization in contrast to fully educated scratchpads. This gives significant length generalization on additions (from 10 to 20 or from 4 to 26 depending on the method) and parities (from 30 to 50-55). For instance, using different methods, [17] can length generalize from 10 to 13 digits for additions, and [11] can get roughly 10 extra bits for parities with moderate accuracy.
@@ 2 Results on the global reasoning barrier
Prior literature. Much work in the learning literature has been devoted to obtaining complexity measures for the sample/time complexity of learning. The largest portion is devoted to target classes in PAC settings, e.g., with the VC dimension measures [18], and some to statistical query (SQ) settings with the statistical dimension measures [14, 19]. Here, we are however interested in measures that are relevant to (1) regular Transformers (or related models) trained by (S)GD, and (2) data distribution fixed by a task. Some recent literature has studied complexity measures for (S)GD-trained neural networks. Various settings and measures have been used, such as the noise sensitivity [20, 6, 21], the cross-predictability [12, 15], the NTK alignment [22, 23], the INAL [24], the $G$-alignment [13], the information and generative exponents [25, 26, 27] and the leap [28]; we refer to Appendix A.2 for discussions on these.
@@ 2.1 Defining globality and auto-regressive globality
The globality degree, or simply globality, of a data distribution measures the least number of input tokens to attend to in order to correlate non-trivially with the label when also given the histogram of the tokens. The specific choice of the mutual information in the definition is not crucial, but one must use a proper measure of dependency (i.e., not just linear correlations), and the mutual information can have convenient chain rule properties. The definition can be related to correlations with $NC^{0}$ circuits (besides for the histogram requirement, see Lemma 6) and also to low-degree polynomial testing, except that it is more general than the latter as it applies to arbitrary token space (without requiring polynomial definitions). Finally, we require the globality to achieve an inverse-polynomial mutual information, the weakest form relevant to weak learning with an inverse-polynomial edge, but one may naturally define the stronger notion with a mutual information of $\Omega(1)$.
@@ 2.2 Transformers require low globality: formal results
(i) In [30], the expressivity of Transformers with constant alphabet size and values recorded to inverse-polynomial accuracy is investigated. It is shown that such a Transformer of constant depth was limited to computing functions in $TC^{0}$. Conversely, it also showed that for any $TC^{0}$ function there is a constant-depth Transformer and instruction string such that when the Transformer is given the instruction string and $x$ as its input it computes the function on $x$. The same technique would extend to show that with logarithmic depth, one can reach $TC^{1}$ (which includes connectivity tasks).
(ii) In [16], well-behaved  [^5] They assumed a constant alphabet size, inverse-polynomial accuracy, constant depth, causal masking, averaging hard attention, and a projected pre-norm. We suspect that all of these requirements other than the constant alphabet size and constant depth could be dropped without changing the results we state here. Transformers with a scratchpad are considered. It is shown that a Transformer with a scratchpad of logarithmic length is limited to computing functions in logspace. We tighten this to $TC^{0}$ in Lemma 5. On the other hand, [16] also shows that any function in $P$ is computable by a Transformer with a scratchpad of polynomial length.
@@
A distribution $P_{X,Y}$ with well-behaved [^6] I.e. $P_{X}$ over $\mathcal{A}^{n}$ such that $|\mathcal{A}|=O(n^{c})$ such that if $X\sim P_{X}$ then there is no value that $X$ takes with probability $O(n^{-c_{1}})$ for any $c_{1}>0$, and every eigenvalue of $X$’s covariance matrix of the indicator functions for the values of $X$’s entries has absolute value $O(n^{-c_{2}}\sum_{i=1}^{n}\mathrm{Var}(X_{i}))$ for some $c_{2}>0$. In other words, the first requirement means that there is no value of $X$ that is frequent enough to weakly learn the function simply by memorizing its value on that input. The second rules out the possibility of an input distribution with a few important components that are easy to notice and determine the output, such as a scenario where there are $\log(n)$ blocks of $n/(2\log(n))$ bits such that all of the bits in a block are always the same and the output depends only on the values of the blocks. Other than that, many distributions of interest are well-behaved, such as the i.i.d. measure as in [12, 25, 24, 28, 26, 27], or measures with dependencies as for the cycle task considered here. $P_{X}$ is efficiently weakly [^7] I.e., with an accuracy that improves on the trivial accuracy by at least $n^{-c}$ for a constant $c$. learnable by a T-regular Transformer if and only if [^8] More specifically, for the converse statement assume that $|\mathcal{A}|$ is constant. $P_{X,Y}$ has constant globality.
@@
(1) An important property of the learning model for the above conjecture is that the probability distribution of the function computed by the model is invariant under permutations of the inputs, and if it is trained in a reasonable way on samples drawn from a probability distribution drawn from a class that is symmetric under permutations of the inputs, its probability distribution will retain its symmetry under permutations of the inputs. For MLPs, we expect most of the results in this paper to apply, with the modification of the globality not having access to the empirical measure of $X$, since one has additional symmetry obtained by exchanging tokens. (2) If we were to use causal masking or a specific choice of positional embeddings that would essentially make it easier for the Transformer to focus on specific subsets of the inputs, and if such subsets are relevant, which may be engineered using some prior knowledge of the target, one could potentially learn functions with higher globality. For instance, we would expect to be unable to learn a function that computes the parity of some subset of $\log(n)$ input bits. However, if we had a positional embedding that gave one value for all of the active bits and mapped all of the other bits to 00, then the Transformer would probably be able to learn the function in question. Likewise, causal masking makes it so that the first few elements of any list the Transformer computes depend only on the first few input symbols, which makes it easier to learn functions that would rely on those symbols.
@@
Let $G$ be a directed graph which consists of a cycle of length $3n$ with probability $2/3$ and 3 cycles of length $n$ otherwise. Next, if there are $3$ cycles pick one vertex from each and if there is one cycle pick 3 vertices that are each $n$ edges apart. Then, label uniformly at random these vertices with $a\_0$, $b\_0$, $c\_0$. Next, number every other vertex by the distance from one of these three to it, and for each $i$, label uniformly at random the vertices at distance $i$ by $a\_i$, $b\_i$, and $c\_i$ and store in $X$ the edges between $a\_{i-1},b\_{i-1},c\_{i-1}$ and $a\_i,b\_i,c\_i$ as described in Figure 2. Finally, let $Y$ report whether $a\_0,b\_0,c\_0$ are in the same cycle or not. Then if we train a T-regular neural network on $(X,Y)$ generated in this manner using population [^9] This would also be true for batch GD with batches of size $n^{c}$, $c$ dependent on the other hyperparameters. gradient descent with polynomial hyperparameters [^10] I.e., either polynomial learning rate, polynomial clipping [12, 31], and weights stored using a logarithmic number of bits of precision and random rounding: for $a<b<c$ if $b$ needs to be rounded to $a$ or $c$ then it rounds to $c$ with probability $(b-a)/(c-a)$, or with polynomial learning rate, polynomial clipping and polynomial noise added to the gradients. and a differentiable loss function then the network fails to weakly learn.
Figure 2: The cycle task variant used in Theorem 1: the above example is stored using the sequence of tokens a_0>b_1;b_0>c_1;c_0>a_1;a_1>a_2;b_1>c_2;c_1>b_2;a_2>b_3;b_2>c_3;c_2>a_3;a_3>b_0;b_3>a_0;c_3>c_0;a_0?b_0?c_0
@@ 3.1 Educated scratchpad
Consider learning parity function $y=f(x_{1},\ldots,x_{n})=x_{1}x_{2}\cdots x_{k}$ where $x_{1},\ldots,x_{n}$ are drawn i.i.d. in {±$\{\pm 1\}$ with uniform distribution. For $k\leq\frac{n}{2}$, one can easily check that the globality of this task is $k$ as any $k-1$ coordinates are independent of the output and the histogram of the tokens does not help. Parity functions are known to be hard to learn [12]. More specifically, it has been previously shown that as $k$ increases the parity task becomes harder to learn to the point that parity of degree $k\wedge(n-k)=\omega(1)$ cannot be learned in $poly(n)$ time with standard poly-size neural networks under standard training assumptions [12]. Note that this is consistent with our results, as the globality is non-constant.
@@
If one has full knowledge of the target function, one could break the target into sub-targets using an educated scratchpad to keep the globality low and thus learn more efficiently (of course one does not have to learn the target under full target knowledge, but one may still want to let a model learn it to develop useful representations in a broader/meta-learning context). One could in theory push this to learning any target that is poly-time computable by emulating a Turing machine in the steps of the scratchpad to keep the overall globality low. Some works have derived results in that direction, such as [33] for some type of linear autoregressive models, or [12] for more abstract neural nets that emulate any Turing machine with SGD training. However, these are mostly theory-oriented works. In practice, one may be instead interested in devising a more ‘generic’ scratchpad. In particular, a relevant feature in many reasoning tasks is the power of induction. This takes place for instance in the last 2 examples (parity and cycle task) where it appears useful to learn inductive steps when possible.

### v3

@@ Abstract
Can Transformers predict new syllogisms by composing established ones? More generally, what type of targets can be learned by such models from scratch? Recent works show that Transformers can be Turing-complete in terms of expressivity, but this does not address the learnability objective. This paper puts forward the notion of globality degree of a target distribution to capture when weak learning is efficiently achievable by regular Transformers. This measure shows a contrast with the expressivity results of Transformers captured by $TC^{0}/TC^{1}$ classes (further studied here), since the globality relates to correlations with the more limited $NC^{0}$ class. We show here experimentally and theoretically under additional assumptions that distributions with high globality cannot be learned efficiently. In particular, syllogisms cannot be composed on long chains. Further, we develop scratchpad techniques and show that: (i) agnostic scratchpads cannot break the globality barrier, (ii) educated scratchpads can break the globality with intermediate steps, although not all such scratchpads can generalize out-of-distribution (OOD), (iii) a notion of ‘inductive scratchpad’, that composes the prior information more efficiently, can both break the globality barrier and improve the OOD generalization. In particular, some of our inductive scratchpads can achieve length generalizations of up to 6×6$\times$6 × for some arithmetic tasks depending on the input formatting.
@@ 1.1 Syllogisms composition
This type of task is nontrivial for current LLMs, and we refer to Appendix I for experiments with GPT models. [^2] At the time of the experiments, ChatGPT was in particular not successful at these two tasks. Note that here we are not interested specifically in solving a graph-based task, but rather in understanding when Transformers can compose and more generally how far they can do so. We would like to identify particular measures on the data distribution (e.g., syllogisms topologies in the above example) that capture when Transformers can efficiently learn.
@@ 1.2 Hardness of long compositions
The input of this task is the graph edge set with the queried vertices. The label is 0 if the two queried vertices are not connected (Class 1) and 1 if they are (Class 2). See Figure 1(a) for an illustration.
@@
(b) Nb. of iterations with $512$ fresh samples to learn ($\geq 95\%$ acc.) the cycle task of size $n$ using GPT2-style models with $10M,25M,85M$ parameters.
@@
A challenge for the cycle task is that there is no clear ‘low-complexity pattern’ in the input representation that indicates whether there are 1 or 2 cycles. No simple statistics based on degrees, edge counts, or finite motif counts that can tell if the vertices are connected or not. One has to consider at least $n$ edges in order to get any correlation with the presence of a path. In other words, the task requires a ‘global reasoning’ involving a ‘large’ number of input tokens and this seems hard for Transformers.
@@ 1.3 Hardness of global reasoning
As discussed, the cycle task appears to be challenging for Transformers as it requires some global reasoning. Other tasks such as subset parities exhibit the same challenge. However the latter can be proved to be not efficiently learnable by various regular neural networks and noisy gradient descent, as one can get explicitly a class of functions (through orbit arguments [12, 13]) that has large statistical dimension [14] or low cross-predictability [12, 15] (see Appendix A.4). For the cycle task, we have a single distribution, and it is unclear how to use the invariances of Transformers to get arguments as in [12, 13], as the input distribution is not invariant under the symmetries of the model. We thus would like to develop a more general complexity measure that unifies why such tasks are hard for Transformer-like models and that formalizes the notion of ‘global reasoning barrier’ when models are trained from scratch. We also would like to understand how the scratchpad methodologies that have proved helpful in various settings (see Section 3) can help here. This raises the questions:
@@ 1.4 Our contributions
    • Theorem 1 that proves the negative side of the above conjecture, the globality barrier, in an instance of the cycle task under certain technical assumptions. (The cycle task is also put forward in the paper as a simple benchmark to test the global reasoning capabilities of models.)
@@
    • Agnostic scratchpad: we extend Theorem 1 to cases where a poly-size scratchpad is used by the Transformers, without any direct supervision of the scratchpad (i.e., the scratchpad mainly provides additional memory/compute). This shows that efficient weak learning is still not possible with such an agnostic scratchpad if the globality is non-constant. An educated guess about what to learn in the scratchpad based on some target knowledge is thus required.
@@
    • We introduce the notion of inductive scratchpad, a type of educated scratchpad that exploits ‘induction’ compared to a fully educated scratchpad and thus composes more efficiently the prior state information. We show that when the target admits an inductive decomposition, such as for the cycle, arithmetic, or parity tasks, the inductive scratchpad both breaks the globality and improves the OOD generalization in contrast to fully educated scratchpads. This gives significant length generalization on additions (from 10 to 18 or from 4 to 26 depending on the method) and parities (from 30 to 50-55). For instance, using different methods, [17] can length generalize from 10 to 12 digits for additions, and [11] can get roughly 10 extra bits for parities.
@@ 2 Results on the global reasoning barrier
Prior literature. Much work in the literature has been devoted to complexity measures for the sample/time complexity of learning. The largest portion is devoted to target classes in PAC settings, e.g., with the VC dimension measures [18], and some to statistical query (SQ) settings with the statistical dimension measures [14, 19]. Here, we are however interested in measures that are relevant to (1) regular Transformers trained by (S)GD, and (2) data distribution fixed by a task. Some recent works have studied complexity measures for (S)GD-trained neural networks. Various settings and measures have been used, such as the noise sensitivity [20, 6, 21], the cross-predictability [12, 15], the NTK alignment [22, 23], the INAL [24], the $G$-alignment [13], the information and generative exponents [25, 26, 27] and the leap [28]; we refer to Appendix A.4 for discussions on these.
@@ 2.1 Defining globality and auto-regressive globality
The globality degree, or simply globality, of a distribution measures the least number of input tokens to attend to in order to correlate non-trivially with the label when also given the histogram of tokens. The specific choice of the mutual information is not crucial, but one must use a proper measure of dependency (i.e., not just linear correlations), and the mutual information can have convenient chain rule properties. The definition can be related to correlations with $NC^{0}$ circuits (besides for the histogram requirement, see Lemma 6) and also to low-degree polynomial testing, except that it is more general than the latter as it applies to arbitrary token space (without requiring polynomial definitions). Finally, we require the globality to achieve an inverse-polynomial mutual information, the weakest form relevant to weak learning with an inverse-polynomial edge, but one may naturally define the stronger notion with a mutual information of $\Omega(1)$.
@@ 2.2 Transformers require low globality: formal results
(i) In [16], the expressivity of Transformers with constant alphabet size and values recorded to inverse-polynomial accuracy is investigated. It is shown that such a Transformer of constant depth was limited to computing functions in $TC^{0}$. Conversely, it also showed that for any $TC^{0}$ function, there is a constant-depth Transformer and instruction string such that when the Transformer is given the instruction string and $x$ as its input it computes the function on $x$. The same technique would extend to show that with logarithmic depth, one can reach $TC^{1}$ (which includes connectivity tasks).
(ii) In [30], well-behaved [^5] They assumed a constant alphabet size, inverse-polynomial accuracy, constant depth, causal masking, averaging hard attention, and a projected pre-norm. We suspect that all of these requirements other than the constant alphabet size and constant depth could be dropped without changing the results we state here. Transformers with a scratchpad are considered. It is shown that a Transformer with a scratchpad of logarithmic length is limited to computing functions in logspace. We tighten this to $TC^{0}$ in Lemma 5 in Appendix H. On the other hand, [30] also shows that any function in $P$ is computable by a Transformer with a scratchpad of polynomial length.
(iii) If we allow Transformers of poly depth then we can convert any poly-sized circuit to a Transformer by replacing each gate in the circuit with an attention head that attends to the values of the appropriate input tokens or attention heads and performs the appropriate computation on them. That means any function in P (including the cycle task) is computable by a Transformer of poly depth and size.
@@
A distribution $P_{X,Y}$ with well-behaved [^6] Conditions such as assuming that there is no value of $X$ that is frequent enough that the model weakly learns the function simply by memorizing the value of that input. Most distributions of interest are well-behaved (including the cycle task). See Definition 6 in Appendix D for a formal definition and additional discussion. $P_{X}$ is efficiently weakly [^7] I.e., with an accuracy that improves on the trivial accuracy by at least $n^{-c}$ for a constant $c$.learnable by a T-regular Transformer if and only if [^8] More specifically, for the converse statement (learning due to constant globality) assume that $|\mathcal{A}|$ is constant.$P_{X,Y}$ has constant globality.
@@
(1) An essential property of the model for the above conjecture is that the probability distribution of the function computed by the model is invariant under permutations of the inputs, and if it is trained reasonably on samples drawn from a distribution drawn from a class that is symmetric under permutations of the inputs, its distribution will retain its symmetry under permutations of the inputs. For MLPs, we expect most of the results in this paper to apply, with the modification of the globality not having access to the empirical measure of $X$, since one has additional symmetry obtained by exchanging tokens. (2) If we were to use causal masking or a specific choice of positional embeddings that would make it easier for the Transformer to focus on specific relevant subsets of the inputs, one could potentially learn functions with higher globality. For instance, we would expect to be unable to learn a function that computes the parity of some subset of $\log(n)$ input bits. However, if we had a positional embedding that gave one value for all of the active bits and mapped other bits to 00, then the Transformer would probably be able to learn the function in question. Likewise, causal masking makes it so that the first few elements of any list the Transformer computes depend only on the first few tokens, which makes it easier to learn functions that would rely on those symbols.
We prove the negative side of Conjecture 1 for a variant of the cycle task.
@@
Let $G$ be a directed graph which consists of a cycle of length $3n$ with probability $2/3$ and 3 cycles of length $n$ otherwise. Next, if there are $3$ cycles pick one vertex from each and if there is one cycle pick 3 vertices that are each $n$ edges apart. Then, label these vertices with $a\_0$, $b\_0$, $c\_0$ uniformly at random. Next, number every other vertex by the distance from one of these three to it, and for each $i$, label uniformly at random the vertices at distance $i$ by $a\_i$, $b\_i$, and $c\_i$. Finally, store the edges between $a\_{i-1},b\_{i-1},c\_{i-1}$ and $a\_i,b\_i,c\_i$ in $X$ (as described in Figure 2), and let $Y$ report whether $a\_0,b\_0,c\_0$ are in the same cycle or not. Then if we train a T-regular neural network on $(X,Y)$ generated in this manner using population [^9] This would also be true for batch GD with batches of size $n^{c}$, $c$ dependent on the other hyperparameters. gradient descent with polynomial hyperparameters [^10] I.e., either polynomial learning rate, polynomial clipping [12, 31], and weights stored using a logarithmic number of bits of precision and random rounding: for $a<b<c$ if $b$ needs to be rounded to $a$ or $c$ then it rounds to $c$ with probability $(b-a)/(c-a)$, or with polynomial learning rate, polynomial clipping and polynomial noise added to the gradients. and a differentiable loss function then the network fails to weakly learn.
The proof of Theorem 1 is presented in Appendix F.
Figure 2: The cycle task variant used in Theorem 1: the above example is stored as a_0>b_1;b_0>c_1;c_0>a_1;a_1>a_2;b_1>c_2;c_1>b_2;a_2>b_3;b_2>c_3;c_2>a_3;a_3>b_0;b_3>a_0;c_3>c_0;a_0?b_0?c_0
@@ 2.3 Agnostic scratchpads cannot break the globality
Next, we put forward a conjecture that agnostic scratchpads (scratchpads without direct supervision on the scratchpad tokens) cannot break the globality barrier. See Appendix E for further discussion.
@@ 3.1 Educated scratchpad
Consider learning parity function $y=f(x_{1},\ldots,x_{n})=x_{1}x_{2}\cdots x_{k}$ where $x_{1},\ldots,x_{n}$ are drawn i.i.d. in {±$\{\pm 1\}$ with uniform distribution. For $k\leq\frac{n}{2}$, one can easily check that the globality of this task is $k$ as any $k-1$ coordinates are independent of the output and the histogram of the tokens does not help. Parity functions are known to be hard to learn [12]. More specifically, it has been previously shown that as $k$ increases the parity task becomes harder to learn to the point that parity of degree $\min\{k,n-k\}=\omega(1)$ cannot be learned in $poly(n)$ time with standard poly-size neural networks under standard training assumptions [12]. Note that this is consistent with our results, as the globality is non-constant.
@@
If one has full knowledge of the target function, one could break the target into sub-targets using an educated scratchpad to keep the globality low and thus learn more efficiently (of course one does not have to learn the target under full target knowledge, but one may still want to let a model learn it to develop useful representations in a broader/meta-learning context). One could in theory push this to learning any target that is poly-time computable by emulating a Turing machine in the steps of the scratchpad to keep the overall globality low. Some works have derived results in that direction, such as [33] for some type of linear autoregressive models, or [12] for more abstract neural nets that emulate any Turing machine with SGD training. However, these are mostly theory-oriented works. In practice, one may be instead interested in devising a more ‘generic’ scratchpad. In particular, a relevant feature in many reasoning tasks is the power of induction. For instance, the parity and cycle tasks are two examples where learning an induction step function appears useful.